# Perfect Wave Transfer in Continuous Quantum Systems

### v1

@@ Abstract
Thus, for a spectrum satisfying the PWT requirements, $\langle\Omega|\varphi(x,T)\varphi(x^{\prime},0)|\Omega\rangle=\langle\Omega|\varphi(-x,0)\varphi(x^{\prime},0)|\Omega\rangle$, meaning that, for fixed $x^{\prime}$, the correlations at $t=T$ are reflections of those at $t=0$. In Fig. 1, we plot Eq. (17) for $K(x)=1$, which, by bosonization, is related to $\log F(t)$ in Eq. (5), in that it encodes the same information; see the SM [60].
Inverse Sturm-Liouville problem.—Reversing the logic, the PWT requirements on the spectrum yield an inverse spectral problem for SL operators with Neumann BCs [and $q(x)=0$]:
Given a spectrum $E_{n}=\pi m_{n}/T$ for integers $m_{n}$ as above, determine the even functions $v(x)$ and $K(x)$.
A related prominent problem is the inverse scattering occurring in the theory of integrable PDEs, whose solution is given by the Gelfand-Levitan-Marchenko equation. However, that case is formulated on an infinite line, while here we are interested in a finite interval.
Before a general discussion, we consider two examples:
The inverse SL problem has a known solution when the spectrum is exactly linear: For $K(x)=K>0$ constant but $v(x)>0$ general, the SL operator is $\mathcal{A}=-v(x)\partial_{x}v(x)\partial_{x}=-v_{0}^{2}\partial_{y}^{2}$ using Eq. (7) with eigenenergies $E_{n}=\pi nv_{0}/L$ and eigenfunctions $u_{n>0}(x)=\sqrt{2v_{0}/L}\cos\bigl{(}n\pi[L+2y]/2L\bigr{)}$ and $u_{0}(x)=\sqrt{v_{0}/L}$. This is precisely the conformal case of non-interacting right/left movers, enjoying PWT for any even $v(x)>0$.
In the previous example, the SL operator was regular. Relaxing this condition by taking $v(x)=v\sqrt{1-(2x/L)^{2}}$ and $K(x)=K\bigl{[}1-(2x/L)^{2}\bigr{]}^{\alpha}$ for $\alpha>-1/2$ and constants $v,K>0$, we get an irregular problem for which Properties (i)(iv) remain valid; the BCs being that a solution is sufficiently well-behaved for Eq. (9) to hold. The eigenfunctions are given by Gegenbauer polynomials $C^{(\alpha)}_{n}(2x/L)$ [80, §18.3] up to normalizations, which have definite parity and satisfy Eq. (13) with $\lambda_{n}=(\pi v_{0}/L)^{2}n(n+2\alpha)$ for $v_{0}=2v/\pi$. Clearly, only the case $\alpha=0$ (Chebyshev polynomials of the first kind) exhibits PWT, while all $\alpha\neq 0$ do not, including Legendre polynomials ($\alpha=1/2$) and Chebyshev polynomials of the second kind ($\alpha=1$) [80, §18.7]. The respective $\varphi\varphi$ correlations are plotted in Fig. 2.
To solve the inverse SL problem in general, we write the operator (11) in Liouville normal form, as is always possible for regular problems [80, §1.13(viii)]: Let $\mathfrak{u}(y)\equiv u(x)\sqrt{K(x)}$ for $y=y(x)$ given by Eq. (7) and define
Figure 2: Plot of $C(x,t)\equiv\operatorname{Re}\langle\Omega|\varphi(x,t)\varphi(-3L/8,0)|\Omega\rangle$ in Eq. (17) for eigenfunctions in the form of (a) Chebyshev polynomials of the first kind, which exhibit PWT, contrasted to (b) Legendre polynomials and (c) Chebyshev polynomials of the second kind, which do not. The upper row contains intensity plots of the correlations as functions of $x/L\in[-1/2,1/2]$ and $t/T\in[0,1]$, while the lower row compares their profiles at $t=0$ and $T$. Note the asymmetry in (b) and (c). The velocity $v(x)$ (same in all cases) and the respective Luttinger parameter $K(x)$ are shown in (d).
@@
with unchanged BCs  [^4] For even $v(x)>0$, they are $v_{0}\partial_{y}\mathfrak{u}(y)-v(x)\Lambda(x)\mathfrak{u}(y)=0$ at $y=\pm L/2$ with $\Lambda(x)\equiv\partial_{x}\ln\sqrt{K(x)}$ [53], which imply Neumann BCs if $K^{\prime}(x)\big{|}_{x=\pm L/2}=0$. and if reflection symmetric in $x$, then so in $y$. Clearly, $\mathfrak{u}_{n}(y)$ have the same eigenvalues as the respective $u_{n}(x)$ and similarly form a complete orthonormal basis but with identity weight.
@@
Results in the literature often require two sets of eigenvalues for different BCs to get a unique solution to the inverse problem. However, one set suffices for reflection-symmetric SL operators and certain BCs, including ours. Indeed, [82] and [83, Theorem II] say that an even $\hat{q}(y)$ is uniquely determined by the spectrum for Neumann BCs. Thus, viewing $\hat{q}(y)$ as given and since $q(x)=0$, Eq. (18) becomes a second-order differential equation for $\sqrt{K(x(y))}$. By general uniqueness results [assuming $\hat{q}(y)$ is continuous] and since $K(x)$ must be even, this determines $K(x)$ up to changing coordinates (7) and an overall rescaling (discarding solutions that are not positive).
The results [82, 83] do not reconstruct $\hat{q}(y)$. However, since we know that $K(x)=K$ gives an SL problem with $m_{n}=n$ and $\hat{q}(y)=0$, a constant Luttinger parameter is the unique solution for the inverse problem given an exactly linear spectrum. Beyond this special case, nothing a priori prevents the spectrum from deviating from the leading $n$ behavior as long as the requirements on the $m_{n}$ remain satisfied. This would then give rise to a class of possible bosonic theories that exhibit PWT, where, by uniqueness, $\hat{q}(y)$ is non-zero, thus $K(x)$ non-constant. While not usually analytically tractable, there are constructive approaches to find $\hat{q}(y)$ [84, 85, 86].
Discussion.—We introduced the notion of PWT and showed that it is exhibited by one-particle excitations in inhomogeneous CFT whenever the velocity profile $v(x)>0$ is even. Breaking conformal invariance, by considering inhomogeneous TLLs with a Luttinger-parameter profile $K(x)>0$, we showed that PWT for general excitations amounts to an inverse spectral problem, which we solved. This continuum result for bosons is remarkably similar to earlier results: For spin chains with nearest neighbor hopping, the corresponding class of operators can be represented as Jacobi matrices, and the resulting inverse eigenvalue problem was studied [8, 87]. As in the continuum, two sets of eigenvalues are generally needed for uniqueness [88, 86], while one set suffices for matrices symmetric with respect to the anti-diagonal (corresponding to reflection symmetry) [89, Theorem 3].
By bosonization, our results for inhomogeneous TLLs cover fermionic theories with position-dependent interactions [53]. Indeed, even when $K(x)=K$ is constant, as long as $K\neq 1$, the theory describes an interacting quantum many-body system, albeit at low-energies.
Lastly, we showed that an inhomogeneous TLL with constant $K(x)=K$ enjoys PWT for general excitations whenever $v(x)>0$ is even. Such a theory is an example of an inhomogeneous CFT. It would be interesting to study PWT for any number of excitations in general CFTs, and also extend to systems on the half or full infinite line.
Acknowledgments.
Acknowledgements.
@@
We are thankful to Pawel Caputa and Shinsei Ryu for motivating discussions on related topics. P.M. gratefully acknowledges financial support from the Wenner-Gren Foundations (Grant No. FT2022-0002). M.C. acknowledges financial support from the European Research Council (ERC Grant Agreement No. 818761), VILLUM FONDEN via the QMATH Centre of Excellence (Grant No. 10059) and the Novo Nordisk Foundation (Grant No. NNF20OC0059939 ‘Quantum for Life’), and thanks the NCCR SwissMAP of the Swiss National Science Foundation and the Section of Mathematics at the University of Geneva for their hospitality.
Supplemental Material
@@
This supplemental material has three parts. Part A gives details for two-point correlations in inhomogeneous 1+1D CFT with open (Neumann) boundary conditions (BCs). Part B describes how to diagonalize general inhomogeneous Tomonaga-Luttinger liquids (TLLs) using Sturm-Liouville (SL) theory, leading to a formulation of perfect wave transfer (PWT) as an inverse SL problem, including extension to massive TLLs. Part C contains details for correlation functions in inhomogeneous TLLs.
@@
where $T_{\pm}(x)$ are the right/left-moving components of the stress-energy tensor in light-cone coordinates, whose Fourier transforms yield two commuting copies of the Virasoro algebra. In position space, they satisfy the equal-time commutation relations
@@
and $\bigl{[}T_{\pm}(x),T_{\mp}(x^{\prime})\bigr{]}=0$. Note that, at non-equal time, $T_{\pm}$ should be understood to only depend on the “right/left-moving coordinate” $x^{\mp}$; for now, we only view these as labels, distinguishing dependencies on right/left components.
Besides the $T_{\pm}(x)$ fields, an important class of operators are Virasoro primary fields $\Phi(x^{-},x^{+})$, defined by how they behave under conformal transformations. The latter are given by orientation-preserving circle diffeomorphisms $f_{+}$, $f_{-}$ and primary fields transform as
@@
where ($\Delta_{\Phi}^{+},\Delta_{\Phi}^{-}$) denote the field’s conformal weights. To impose open BCs at $x=\pm L/2$, as in the main text, we will assume that primary fields come in pairs $\Phi_{+}$, $\Phi_{-}$ with conformal weights satisfying $\Delta^{+}_{\Phi_{+}}=\Delta^{-}_{\Phi_{-}}$, $\Delta^{-}_{\Phi_{+}}=\Delta^{+}_{\Phi_{-}}$, since this allows one to couple excitations so that right movers are reflected to left movers when reaching $x=L/2$, and vice-versa at $x=-L/2$. To ease notation, we will use $\Delta^{+}_{\Phi}\equiv\Delta^{+}_{\Phi_{+}}$, $\Delta^{-}_{\Phi}\equiv\Delta^{-}_{\Phi_{+}}$ for each such pair.
@@ 1. Example 1:
    For the inhomogeneous free fermions in the main text, the components of the stress-energy tensor are $T_{\pm}(x)=(1/2)\bigl{[}\!\left.:\!\psi^{\dagger}_{\pm}(x)(\mp\mathrm{i}\partial_{x})\psi^{\vphantom{\dagger}}_{\pm}(x)\!:\right.+\mathrm{h.c.}\bigr{]}-\pi\big{/}12L^{2}$, and the primary fields of the theory are the fermionic fields $\psi_{\pm}(x)$ themselves, whose conformal weights are $\Delta^{+}_{\psi_{+}}=\Delta^{-}_{\psi_{-}}=1/2$, $\Delta^{-}_{\psi_{+}}=\Delta^{+}_{\psi_{-}}=0$.
@@ 2. Example 2:
    For an inhomogeneous TLL with constant Luttinger parameter $K(x)=K$, one has $T_{\pm}(x)=({1}/{4\pi K})\!\left.:\!\bigl{[}\pi\Pi(x)\mp K\partial_{x}\varphi(x)\bigr{]}^{2}\!:\right.-\pi\big{/}12L^{2}$, and the primary fields include vertex operators constructed as exponentials of the bosonic fields. Applied to the Luttinger model [46, 47, 48], these vertex operators are (renormalized) fermionic fields $\psi_{\pm}$ with conformal weights $\Delta^{+}_{\psi_{+}}=\Delta^{-}_{\psi_{-}}=(1+K)^{2}/8K$, $\Delta^{-}_{\psi_{+}}=\Delta^{+}_{\psi_{-}}=(1-K)^{2}/8K$. (Clearly, free fermions are recovered for $K=1$.)
@@
for all pairs $\Phi_{+},\Phi_{-}$ of Virasoro primary fields. It follows that we can equivalently consider a chiral inhomogeneous CFT of only right movers on the interval $[-L/2,3L/2]$ of length $2L$ with velocity
$\underline{v}(x)\equiv\begin{cases}v(x)&\text{for }x\in[-L/2,L/2],\\ v(L-x)&\text{for }x\in[L/2,3L/2]\end{cases}$ (24)
@@
for smooth enough functions $\xi^{\pm}_{j}(x)$, interpreted as distributions or “wave packets” of right/left-moving excitations in each state, and some $k\in\mathbb{R}$ in order to accommodate for $k_{\textrm{F}}$ in Eq. (3). (The latter will have no effect on PWT, and one might as well set $k=0$.) As in the main text, this is a restricted class of states which only allow for one-particle excitations, while we eased other restrictions in the sense that the excitations are not necessarily fermionic and the distributions of right/left movers can be different.
@@
in order to eventually compare $|F(t)|$ with $|\langle\Phi_{1}|\Phi_{1}\rangle|$ for specular waves, $\xi^{\pm}_{2}(x)=\xi^{\mp}_{1}(-x)$. Using the unfolding relations (23), we obtain
@@
where we used that $H_{\textrm{iCFT}}|\Omega\rangle=C|\Omega\rangle$ for some real constant $C$ that encodes the Casimir energy. (This only gives rise to an overall phase that one could safely ignore.) Clearly, as stated in the main text, the above amounts to evaluating two-point correlations for primary fields, specifically those of right movers in a chiral CFT on $[-L/2,3L/2]$.
First, as a preliminary step, for primary fields with periodic boundary conditions on the system with length $2L$, conformal invariance implies that the undeformed two-point correlations are
@@
where $G^{\pm}_{\Phi\Phi}(x)$ are $2L$-periodic functions. For example, as can be seen from, e.g., [55], the correlations for right-moving free fermions (cf. Example 1 above) on the system with length $2L$ and periodic BCs are
@@
where we used that $\Delta^{+}_{\psi}=1/2$, $\Delta^{-}_{\psi}=0$. Similarly, for the Luttinger model (cf. Example 2 above), the correlations for the (renormalized) right-moving fermionic fields are
@@
Second, by results spelled out in Proposition 3.1 in [26] for inhomogeneous CFT, the time dependence in $F(t)$ means that those correlations are evaluated at spacetime points given by $\underline{x}^{\pm}_{t}(x)$ in Eq. (27). In addition, the action of $U_{+}(\underline{f})$ from the expression (29) for $|\Omega\rangle$ also leads to a conformal transformation of the fields. It follows from this and the properties (27) that
@@
$F(t)=\mathrm{e}^{-\mathrm{i}Ct}\frac{1}{2\pi}\int_{-L/2}^{L/2}\mathrm{d}x_{1}\,\int_{-L/2}^{L/2}\mathrm{d}x_{2}\,\underline{f}^{\prime}(x_{2})^{\Delta^{+}_{\Phi}}\underline{f}^{\prime}(x_{1})^{\Delta^{+}_{\Phi}}\underline{f}^{\prime}(x_{2})^{\Delta^{-}_{\Phi}}\underline{f}^{\prime}(x_{1})^{\Delta^{-}_{\Phi}}\\ \begin{aligned} \times\biggl{(}&\mathrm{e}^{\mathrm{i}k(x_{2}-x_{1})}\overline{\xi^{+}_{2}(x_{2})}\xi^{+}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}\underline{f}(x_{2})-\underline{f}(x_{1})-\underline{v}_{0}t\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{v}_{0}t\bigr{)}^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{\mathrm{i}k(x_{2}+x_{1})}\overline{\xi^{+}_{2}(x_{2})}\xi^{-}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{f}(L)-\underline{v}_{0}t\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{f}(L)+\underline{v}_{0}t\bigr{)}^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{-\mathrm{i}k(x_{2}+x_{1})}\overline{\xi^{-}_{2}(x_{2})}\xi^{+}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}-\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{f}(L)-\underline{v}_{0}t\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}-\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{f}(L)+\underline{v}_{0}t\bigr{)}^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{-\mathrm{i}k(x_{2}-x_{1})}\overline{\xi^{-}_{2}(x_{2})}\xi^{-}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}-\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{v}_{0}t\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}-\underline{f}(x_{2})+\underline{f}(x_{1})+\underline{v}_{0}t\bigr{)}^{2\Delta^{-}_{\Phi}}\biggr{)},\end{aligned}$ (41)

### v3

@@ Abstract
Thus, for a spectrum satisfying the PWT requirements, $\langle\Omega|\varphi(x,T)\varphi(x^{\prime},0)|\Omega\rangle=\langle\Omega|\varphi(-x,0)\varphi(x^{\prime},0)|\Omega\rangle$, meaning that, for fixed $x^{\prime}$, the correlations at $t=T$ are reflections of those at $t=0$. In Fig. 1, we plot Eq. (17) for $K(x)=1$, which, by bosonization, is related to $\log F(t)$ in Eq. (5), in that it encodes the same information [74].
Inverse Sturm-Liouville problem—Reversing the logic, the spectral requirements for PWT yield an inverse problem for SL operators with Neumann BCs [101]: Given a spectrum $E_{n}=\pi m_{n}/T$ for integers $m_{n}$ as above, determine the even functions $v(x)$, $K(x)$ [and $q(x)$ if nonzero].
Before a general discussion, we consider two examples.
First, the solution is known if the spectrum is exactly linear: For $K(x)=K$ constant but $v(x)$ general, then $\mathcal{A}=-v(x)\partial_{x}v(x)\partial_{x}=-v_{0}^{2}\partial_{y}^{2}$ using Eq. (7) with $E_{n}=\pi nv_{0}/L$ and $u_{n>0}(x)=\sqrt{2v_{0}/L}\cos\bigl(n\pi[L+2y]/2L\bigr)$, $u_{0}(x)=\sqrt{v_{0}/L}$. This is the conformal case of uncoupled right and left movers, enjoying PWT for any even $v(x)>0$.
Second, for $v(x)=v\sqrt{1-(2x/L)^{2}}$ and $K(x)=K\bigl[1-(2x/L)^{2}\bigr]^{\alpha}$ with $\alpha>-1/2$ and $v,K>0$, we get an irregular problem possessing Properties (i)(iv) [102]. The eigenfunctions are Gegenbauer polynomials $C^{(\alpha)}_{n}(2x/L)$ (up to normalizations), which have definite parity and satisfy Eq. (13) with $\lambda_{n}=(\pi v_{0}/L)^{2}n(n+2\alpha)$, $v_{0}=2v/\pi$; cf. Sec. 18.3 in [103]. Clearly, only $\alpha=0$ (Chebyshev polynomials of the first kind) exhibits PWT, while all $\alpha\neq 0$ do not, including $\alpha=1/2$ (Legendre polynomials) and $\alpha=1$ (Chebyshev polynomials of the second kind); cf. Sec. 18.7 in [103]. The respective $\varphi\varphi$ correlations are plotted in Fig. 2.
To solve the inverse SL problem in general, we write the operator (11) in Liouville normal form, as is always possible for regular problems; cf. Sec. 1.13(viii) in [103]: Let $\mathfrak{u}(y)\equiv u(x)\sqrt{K(x)}$ for $y=y(x)$ in Eq. (7) and define
@@
with unchanged BCs [104], and if reflection symmetric in $x$, then so in $y$. Clearly, $\mathfrak{u}_{n}(y)$ have the same eigenvalues as the respective $u_{n}(x)$, and they similarly form a complete orthonormal basis but with identity weight.
@@
Two sets of eigenvalues for different BCs are usually required to uniquely solve an inverse SL problem. However, one set suffices for reflection-symmetric operators with certain BCs. Indeed, an even $\hat{q}(y)$ is uniquely determined by the spectrum for Neumann BCs [105, 106]. Thus, taking $\hat{q}(y)$ as given, Eq. (18) [with $q(x)=0$] becomes a second-order differential equation for $\sqrt{K(x(y))}$. By uniqueness, this determines an even $K(x)$ up to changing coordinates as in Eq. (7) and an overall rescaling [assuming that $\hat{q}(y)$ is continuous and the solution $\sqrt{K(x)}$ is positive].
The results [105, 106] do not reconstruct $\hat{q}(y)$, although there are constructive approaches [107, 108, 109]. However, as $K(x)=K$ gives an SL problem with $m_{n}=n$ and $\hat{q}(y)=0$, a constant Luttinger parameter is the unique solution for the inverse problem given an exactly linear spectrum. Beyond this case, one may imagine that the spectrum could deviate from the leading $n$ behavior as long as the requirements on the $m_{n}$ are satisfied. By an exact Bohr-Sommerfeld/WKB quantization condition [110, 111, 112, 113], we show in the SM [74] that this depends on the regularity: If the system is regular also after unfolding the interval [L/2,L/2][-L/2,L/2] to the circle of twice the length [114], then PWT requires $\hat{q}(y)=0$. Thus, the conformal case of an exactly linear spectrum is the only possibility for such regular inhomogeneous TLLs to exhibit PWT.
Discussion—We investigated the collective transfer of quantum information in continuous communication channels, introducing the notion of perfect wave transfer (PWT). Assuming conformal invariance, we showed that PWT is exhibited for one-particle excitations whenever the system is reflection symmetric. More generally, PWT with respect to arbitrary states amounts to an inverse spectral problem, studied in detail for inhomogeneous TLL theory—a prototypical model of 1+1D bosons with coupled right and left movers. Its solution shows that conformal invariance is essential for PWT if sufficient regularity is required. This approach is similar to results for spin chains with nearest-neighbor hopping: Their Hamiltonians can be represented as Jacobi matrices, yielding an inverse eigenvalue problem [9, 115] whose solution is unique [116, 109] given one set of eigenvalues for matrices symmetric with respect to the antidiagonal (reflection symmetry) [117]. Remarkably, however, the continuum problem is more constrained due to asymptotic requirements, which imply that the spectrum must be exactly linear for inhomogeneous bosons to exhibit PWT under strong enough regularity conditions. This is not the case for discrete systems due to their intrinsic ultraviolet cutoff. Such a cutoff is also relevant for applications of TLL theory [49], and if present, there are no asymptotic requirements and more possibilities for PWT even for regular inhomogeneous TLLs. Moreover, as shown in the SM [74], PWT is also possible in inhomogeneous TLLs with fine-tuned mass terms, governed by a similar inverse spectral problem but with $q(x)\neq 0$.
By bosonization, our results for inhomogeneous TLLs cover fermionic theories with position-dependent interactions [61]. The essential property for these systems to exhibit PWT as in Eq. (1) was shown to be conformal invariance, assuming that the inhomogeneities are sufficiently regular, while it remains unclear what is possible if the regularity is relaxed. Consequently, PWT for all (not only one-particle) states holds for particular CFTs, such as inhomogeneous CFTs of free fermions and compactified bosons. It would be interesting to prove if PWT holds for any number of excitations in general CFTs and to extend to systems on the half or full infinite line.
Our Letter can be seen as a first investigation of PWT for continuous quantum systems, focusing on universal low-energy descriptions within CFT and TLL theory. A second step would be to study PWT for specific gapless 1+1D models, including Lieb-Liniger [85]. For those, our results apply at low energies, while full treatments require tailored solution methods, such as Bethe ansatz, but typically extended to inhomogeneous systems for PWT to be possible—a very difficult problem in itself. Generic interaction or curvature effects may also spoil PWT beyond the low-energy regime. Refinements may then be needed, either relaxing PWT to specific classes of observables or states, or reformulating it as almost PWT with fidelity $1-\varepsilon$ ($\varepsilon>0$) for the transfer of information. A concrete proposal for a general approach that builds on our present results would be to use nonlinear TLL theory [118, 119] to study such questions.
@@
Acknowledgments—We are thankful to Pawel Caputa, Shinsei Ryu, and Imke Schneider for motivating discussions. P.M. acknowledges financial support from the Wenner-Gren Foundations (Grant No. FT2022-0002). M.C. acknowledges financial support from the European Research Council (Grant No. 818761), Villum Fonden via the QMATH Centre of Excellence (Grant No. 10059), and the Novo Nordisk Foundation (Grant No. NNF20OC0059939 “Quantum for Life”) and thanks the NCCR SwissMAP of the Swiss National Science Foundation and the Section of Mathematics at the University of Geneva for their hospitality. We also thank the organizers of the workshop “Integrability in Condensed Matter Physics and Quantum Field Theory” in 2023 at the SwissMAP Research Station in Les Diablerets, where this work was initiated.
Supplemental Material for:
Perfect Wave Transfer in Continuous Quantum Systems
@@
This supplemental material has three parts. Part A gives details for two-point correlations in inhomogeneous 1+1D conformal field theory (CFT) with open boundary conditions (BCs). Part B describes how to diagonalize general inhomogeneous Tomonaga-Luttinger liquids (TLLs) using Sturm-Liouville (SL) theory, leading to the formulation of perfect wave transfer (PWT) as an inverse SL problem, including extension to “massive” TLLs and consequences inferred from Weyl’s law and the Bohr-Sommerfeld/WKB quantization condition. Part C contains details on correlations in inhomogeneous TLLs.
@@
where $T_{\pm}(x)$ are the right- (++) and left- (-) moving components of the stress-energy tensor in light-cone coordinates, whose Fourier transforms yield two commuting copies of the Virasoro algebra; see, e.g., [36]. In position space, they satisfy the equal-time commutation relations
@@
and $\bigl[T_{\pm}(x),T_{\mp}(x^{\prime})\bigr]=0$, where $c$ is the so-called central charge. Note that $T_{\pm}$ should be understood to only depend on the “right/left-moving coordinate” $x^{\mp}$, i.e., $T_{\pm}(x^{\mp})$; for now, we only view these as labels, distinguishing dependencies for right- or left-moving components.
Besides the $T_{\pm}(x)$ fields, an important class of operators are Virasoro primary fields $\Phi(x^{-},x^{+})$, defined by how they behave under conformal transformations. The latter are given by orientation-preserving circle diffeomorphisms $f_{+}$ and $f_{-}$, under which primary fields transform as
@@
where ($\Delta_{\Phi}^{+},\Delta_{\Phi}^{-}$) denote the field’s conformal weights. To impose open BCs at $x=\pm L/2$, as in the main text, we will assume that primary fields come in pairs $\Phi_{+}$, $\Phi_{-}$ with conformal weights satisfying $\Delta^{+}_{\Phi_{+}}=\Delta^{-}_{\Phi_{-}}$, $\Delta^{-}_{\Phi_{+}}=\Delta^{+}_{\Phi_{-}}$, since this allows one to couple excitations so that right movers are reflected to left movers when reaching $x=L/2$, and vice versa at x=L/2x=-L/2. For ease of notation, we will use $\Delta^{+}_{\Phi}\equiv\Delta^{+}_{\Phi_{+}}$, $\Delta^{-}_{\Phi}\equiv\Delta^{-}_{\Phi_{+}}$ for each such pair.
@@ 1. Example 1:
    For the inhomogeneous free fermions in the main text, the components of the stress-energy tensor are $T_{\pm}(x)=(1/2)\bigl[\!\left.:\!\hskip-0.5pt\psi^{\dagger}_{\pm}(x)(\mp\mathrm{i}\partial_{x})\psi^{\vphantom{\dagger}}_{\pm}(x)\hskip-0.5pt\!:\right.+\mathrm{h.c.}\bigr]-\pi\big/12L^{2}$, the central charge is $c=1$, and the primary fields of the theory are the fermionic fields $\psi_{\pm}(x)$ themselves, whose conformal weights are $\Delta^{+}_{\psi_{+}}=\Delta^{-}_{\psi_{-}}=1/2$, $\Delta^{-}_{\psi_{+}}=\Delta^{+}_{\psi_{-}}=0$.
@@ 2. Example 2:
    For an inhomogeneous TLL with constant Luttinger parameter $K(x)=K$, one has $T_{\pm}(x)=({1}/{4\pi K})\!\left.:\!\hskip-0.5pt\bigl[\pi\Pi(x)\mp K\partial_{x}\varphi(x)\bigr]^{2}\hskip-0.5pt\!:\right.-\pi\big/12L^{2}$, the central charge is $c=1$, and the primary fields include vertex operators constructed as exponentials of the bosonic fields. Applied to the Luttinger model [54, 55, 56], these vertex operators are (renormalized) fermionic fields $\psi_{\pm}$ with conformal weights $\Delta^{+}_{\psi_{+}}=\Delta^{-}_{\psi_{-}}=(1+K)^{2}/8K$, $\Delta^{-}_{\psi_{+}}=\Delta^{+}_{\psi_{-}}=(1-K)^{2}/8K$. (Free fermions are recovered by setting $K=1$.)
@@
for all pairs $\Phi_{+},\Phi_{-}$ of Virasoro primary fields. It follows that we can equivalently consider a chiral inhomogeneous CFT of only right movers on the interval [L/2,3L/2][-L/2,3L/2] of length $2L$ with velocity
$\underline{v}(x)\equiv\begin{cases}v(x)&\text{for }x\in[-L/2,L/2],\\ v(L-x)&\text{for }x\in[L/2,3L/2],\end{cases}$ (24)
@@
for smooth enough functions $\xi^{\pm}_{j}(x)$, interpreted as “waves” of right- or left-moving excitations in each state, and some $k\in\mathbb{R}$ in order to accommodate for $k_{\textrm{F}}$ in Eq. (3). (The latter will have no effect on PWT, and one might as well set $k=0$.) As in the main text, this is a restricted class of states which only allow for one-particle excitations, while we eased other restrictions in the sense that the excitations are not necessarily fermionic and the distributions of right and left movers in the “waves” can be different.
@@
in order to eventually compare $|F(t)|$ with $|\langle\Phi_{1}|\Phi_{1}\rangle|$ for specular waves, $\xi^{\pm}_{2}(x)=\xi^{\mp}_{1}(-x)$. Using the unfolding relations in Eq. (23), we obtain
@@
where we used that $H_{\textrm{iCFT}}|\Omega\rangle=C|\Omega\rangle$ for some real constant $C$ that encodes the Casimir energy. (This only gives rise to an overall phase that one could safely ignore.) Clearly, as stated in the main text, the above amounts to evaluating and comparing two-point correlations for primary fields, specifically those of right movers in a chiral CFT on [L/2,3L/2][-L/2,3L/2].
First, as a preliminary step, for primary fields in a system of length $2L$ with periodic BCs, conformal invariance implies that the (undeformed) two-point correlations are of the form
@@
where $G^{\pm}_{\Phi\Phi}(x)$ are $2L$-periodic functions. For example, as can be seen from, e.g., [63, 67], the two-point correlations for right-moving free fermions (cf. Example 1 above) are
@@
where we used that $\Delta^{+}_{\psi}=1/2$, $\Delta^{-}_{\psi}=0$. Similarly, for the Luttinger model (cf. Example 2 above), the two-point correlations for the (renormalized) right-moving fermionic fields are
@@
Second, by results spelled out in Proposition 3.1 in [36] for inhomogeneous CFT, the time dependence in $F(t)$ means that those correlations are evaluated at spacetime points given by $\underline{x}^{\pm}_{t}(x)$ in Eq. (27). In addition, the action of $U_{+}(\underline{f})$ from the expression (29) for $|\Omega\rangle$ also leads to a conformal transformation of the fields. It follows from this and the properties in Eq. (27) that
@@
$\hskip-7.11317ptF(t)=\mathrm{e}^{-\mathrm{i}Ct}\frac{1}{2\pi}\int_{-L/2}^{L/2}\mathrm{d}x_{1}\,\int_{-L/2}^{L/2}\mathrm{d}x_{2}\,\underline{f}^{\prime}(x_{2})^{\Delta^{+}_{\Phi}}\underline{f}^{\prime}(x_{1})^{\Delta^{+}_{\Phi}}\underline{f}^{\prime}(x_{2})^{\Delta^{-}_{\Phi}}\underline{f}^{\prime}(x_{1})^{\Delta^{-}_{\Phi}}\\ \begin{aligned} \times\biggl(&\mathrm{e}^{\mathrm{i}k(x_{2}-x_{1})}\overline{\xi^{+}_{2}(x_{2})}\xi^{+}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl(\underline{f}(x_{2})-\underline{f}(x_{1})-\underline{v}_{0}t\bigr)^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl(\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{v}_{0}t\bigr)^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{\mathrm{i}k(x_{2}+x_{1})}\overline{\xi^{+}_{2}(x_{2})}\xi^{-}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl(\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{f}(L)-\underline{v}_{0}t\bigr)^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl(\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{f}(L)+\underline{v}_{0}t\bigr)^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{-\mathrm{i}k(x_{2}+x_{1})}\overline{\xi^{-}_{2}(x_{2})}\xi^{+}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl(-\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{f}(L)-\underline{v}_{0}t\bigr)^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl(-\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{f}(L)+\underline{v}_{0}t\bigr)^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{-\mathrm{i}k(x_{2}-x_{1})}\overline{\xi^{-}_{2}(x_{2})}\xi^{-}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl(-\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{v}_{0}t\bigr)^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl(-\underline{f}(x_{2})+\underline{f}(x_{1})+\underline{v}_{0}t\bigr)^{2\Delta^{-}_{\Phi}}\biggr),\end{aligned}$ (41)